None of the cars finished the race here is what happend
car A broke 1/9th into last 3rd of race. How far did car A go?
car B completed the first 1/2 of 13/16th of the race how far did car B go?
car C broke with 1/3 left of race how far did car C go?
car D broke with 1/5 of the last 5th of race completed, how far did D go?
car E went 1/2 as far as car B how far did car E go?
Car F went 2/3 as far as car B how far did car f go?
Who went the farthest? who went the shortest.


Answer by longjonsilver(2297)

A did 2/3 + (1/9 of 1/3) = 2/3 + 1/9 x 1/3) = 2/3 + 1/27 = 18/27 + 1/27 = 19/27
B did 1/2x13/16 = 13/32
C did 2/3
D did 4/5 + (1/5 of 1/5) = 4/5 + 1/25 = 20/25 + 1/25 = 21/25
E did 1/2 of 13/32 = 1/2 x 13/32 = 13/64
F did 2/3 of 13/32 = 2/3 x 13/32 = 13/48
Both E and F did less than B, so ignore them. It is between A, B, C and D. Looking at B, this is a small fraction compared to the other 3.
Converting to decimals on calculator (you didn't say i couldn't)...
A = 0.704
C = 0.666
D = 0.84
Answer...D went furthest.
B was shortest of A,B,C and D and then E and F are both fractions of B, so it had to be one of E or F. 1/2 is less than 2/3, so E went shortest distance.
jon